OpenStudy (anonymous):

a bag contains only red, yellow and blue discs. the probability of drawing a red disc at random is 1/2 the probability of drawing a yellow disc at random is x the probability of drawing a blue disc at random is 4x one disc is to be selected at random. work out the probability that it will be a blue disc. give your answer as a numerical value. could someone help me please:)

OpenStudy (anonymous):

i am also a beginner .. i can hint u .. should i ?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

as u told, the probability of drawing red ball is 1/2 it means there are only 2 balls in the bag and one of them is red.. just think over it

OpenStudy (anonymous):

oh wait dont worry i think ive done it :)

OpenStudy (anonymous):

did my answer helped u or not ? :) plz tell me plz tell me i'm also very excited :)

OpenStudy (anonymous):

no not really aha, sorry! but thanks anyway!

OpenStudy (anonymous):

omg.. haahahahahaahahaha.... hahahaaha h.. i'm laughing .. haahahah n thanks too but why u gave me medal :(

OpenStudy (anonymous):

because you helped :) and im a nice person ahah

OpenStudy (anonymous):

you r nice person..... let me think... yah u r nice person bcoz you gave me medal although my answer didn't help u :)

OpenStudy (anonymous):

if u know the answer tell me how to do that .. bcoz i'm also doing probability it will help me too .. i helped u na now u help me

OpenStudy (anonymous):

prob. of red disc+prob/ of yellor disc + prob. of blue disc=prob of getting a disc 1/2+x+4x=1 1/2+5x=1 5x=1/2 x=1/10 so the prob. of getting blue disc is... 4x=4(1/10)=2/5 < answer

OpenStudy (anonymous):

well i just gave the amoutn of discs a random number, say 10 and because half of them must be red that gives: 10-5=5 then because there is: yellow=x blue=4x you must put it into a ratio: yellow 1:4 blue yellow= 1/10 blue= 4/10 answer is 4/10 :)

OpenStudy (anonymous):

yes and upon simplyfying you'll have 2/5

OpenStudy (anonymous):

ohhh... thanks alot .. :) :)

OpenStudy (anonymous):

i gave u medal for gud question now account is finished .. :)